Logical signals for the Left PMTs of S2m had a wrong map for the channels 9-16. For example PMT #9 was connected to channel #16. Connections for HV and analog channels were correct. Connections of logical channels for the Left PMts 1-8 were correct.
I fixed all problems and presently connections are correct.